# Scalarisation-based risk concepts for robust multi-objective optimisation

### v1

@@ 2.3 Risk functionals
Most existing work in robust multi-objective optimisation has focussed on the simple scenario where each uncertain parameter $\boldsymbol{\xi}\in\Xi$ is given equal weight. In this work, we consider a more general treatment where the uncertain parameter is treated as a random variable that is defined sensibly over the standard probability space. Precisely, we suppose that the uncertain parameter $\boldsymbol{\xi}\in\Xi$ is distributed according to some probability density $p\in\mathcal{P}[\Xi]$ that could potentially be known but might in general be unknown. Equipped with this set-up, the value of an input $\mathbf{x}\in\mathbb{X}$ will no longer be determined by its equally weighted output set (7), but instead will be defined using a sensible weighted alternative. More specifically, we will consider the use of risk functionals (Rüschendorf, 2013, Part 2) in order to obtain these proxy set of values. For this reason, we will now give a brief overview on the concepts of a univariate risk functional (Section 2.3.1) and a multivariate risk functional (Section 2.3.2). In Figure 2, we give an illustration of some univariate and multivariate risk functionals for a simple one-dimensional and two-dimensional example, respectively.
@@ 2.3.1 Univariate risk functionals
In the following sections, we will consider the problem of maximising the output of a risk functional. As a result, we will often refer to the resulting output of a univariate risk functional as the risk-adjusted or risk-penalised objective value. Naturally, we want these risk functional to satisfy some sensible properties in order for the corresponding risk-penalised optimisation problem to make sense from a modelling perspective. For this reason, we focus our attention on partially coherent univariate risk functional (Definition 2.4). By design, these functionals satisfy four additional properties which are sufficient enough to prove many of the results which we state and describe later in Section 4. Note that all of the univariate risk functionals listed above (Examples 2.2, 2.3, 2.4 and 2.5) are partially coherent.
@@ 2.3.2 Multivariate risk functionals
$\boldsymbol{\rho}^{\textnormal{WorstCase}}[h]$ :=$:=\{(\inf_{\boldsymbol{\xi}\in\mathcal{E}}h^{(1)}(\boldsymbol{\xi}),\dots,\inf_{\boldsymbol{\xi}\in\mathcal{E}}h^{(M)}(\boldsymbol{\xi}))\in\mathbb{R}^{M}\},$
$\boldsymbol{\rho}^{\textnormal{BestCase}}[h]$ :=$:=\{(\sup_{\boldsymbol{\xi}\in\mathcal{E}}h^{(1)}(\boldsymbol{\xi}),\dots,\sup_{\boldsymbol{\xi}\in\mathcal{E}}h^{(M)}(\boldsymbol{\xi}))\in\mathbb{R}^{M}\}.$
@@
Given a probability density $p\in\mathcal{P}[\Xi]$, the multivariate expectation functional maps any vector-valued function $h:\Xi\rightarrow\mathbb{R}^{M}$ to the singleton set comprised of its corresponding expectation
$\boldsymbol{\rho}^{\textnormal{Exp}}[h]:=\{\mathbb{E}_{p(\boldsymbol{\xi})}[h(\boldsymbol{\xi})]\in\mathbb{R}^{M}\}.$
For example, Deb and Gupta (2005) considered a special case where the uncertain parameter models the input noise in a standard vector-valued optimisation problem (2), that is $\boldsymbol{\rho}^{\textnormal{Exp}}[f(\mathbf{x},\cdot)]=\{\mathbb{E}_{p(\boldsymbol{\xi})}[g(\mathbf{x}+\boldsymbol{\xi})]\in\mathbb{R}^{M}\}$ for any input $\mathbf{x}\in\mathbb{X}$, where $p\in\mathcal{P}[\Xi]$ is the input noise distribution and $g:\mathbb{X}\rightarrow\mathbb{R}^{M}$ is the vector-valued objective function.
@@
$\boldsymbol{\rho}^{\textnormal{CW}}[h]:=\{(\rho^{(1)}[h^{(1)}],\dots,\rho^{(M)}[h^{(M)}])\in\mathbb{R}^{M}\},$
@@ 3 Robust multi-objective optimisation
Several notions of robustness have already been proposed for the robust multi-objective optimisation problem (1). Regardless of these different definitions, the most prominent solution strategies that have been proposed to solve these robust problems are all largely based on the scalarisation methodology. We will now begin this section with Section 3.1, where we give a more precise recap on some of the prominent ideas and strategies that have been proposed in these earlier work. In particular, our brief review is designed to highlight the interplay between the two core computational routines that appear in all of these existing approaches: robustification and scalarisation. Inspired by these connections, we move on to Section 3.2 and Section 3.3, where we introduce the two main solution strategies that will be the primary focus of this work: the robustify then scalarise (12) approach and the scalarise then robustify (15) approach. Notably, these two family of solution strategies encapsulate all of the existing strategies described in Section 3.1 as special cases. Lastly, we end this section with Section 3.4, where we present a discussion that attempts to answer a number of sensible questions regarding these two computationally-driven strategies.
3.1 Robust partial ordering
@@
In general, we cannot solve the set-valued optimisation problem (8) in its entirety. Instead, most existing optimisation strategies work by identifying some finite approximation to this problem based on some sensible adaptation to the scalarisation approach (Section 2.2). To elaborate, one usually aims to define and solve a finite collection of scalar-valued optimisation problems whose solutions are robust with respect to the set partial ordering of interest. Below, we will recall some of these examples for the upper and lower set ordered problem. Moreover, we highlight how these examples are just special cases of our RTS and STR approach, which we will define later in Section 3.2 and Section 3.3, respectively.
@@
3.2 Robustify then scalarise
Union robust optimisation problem
Recall from Section 2.2 that it is possible to recover any weakly Pareto optimal solution by solving a collection of length scalarised problems—see Theorem 2.1. We will now present the corresponding analogue of this result for the RTS setting. To elaborate, our main result in eq. 14 shows that the corresponding RTS length scalarised problem (14) can be used in order to target all of the weak solutions to the bounded vector-valued optimisation problem
Proposition 3.1 (Union robust weak solutions)
3.3 Scalarise then robustify
STR length scalarised problem
3.4 Discussion
The RTS and STR approach are two general computationally-driven strategies for robust multi-objective optimisation. These two approaches are very flexible and lightweight because they only require the decision maker to specify a sensible family of scalarisation functions and a risk functional in order to define robustness. The former set of functions is determined by the decision maker’s internal preferences and specific goals. Whilst the latter functional is determined by the uncertainties in the problem and the decision maker’s adversity to risk. In this section, we look at some natural questions that one might have regarding these two approaches, such as: which approach should one use in practice (Section 3.4.1) and when are these two approaches equivalent (Section 3.4.2)? Moreover, as an aside, we also relate some of these ideas with existing concepts from multi-objective reinforcement learning (Section 3.4.3).
@@ 3.4.1 When to robustify and scalarise?
In practice, decision makers are interested in making principled decisions that can be justified and explained to other stakeholders. From an interpretation perspective, the RTS approach is appealing because it can easily be interpreted as a standard multi-objective optimisation problem (2) wherein one replaces the objective function with a risk-penalised version (13). The corresponding scalarisation methodology can then be viewed as an operational technique in order to acquire the robust points which are the most desirable. In contrast, the STR approach takes a much more preference-driven strategy that is based largely on the scalarisation methodology. Specifically, one has to first identify a scalarisation function that describes the most preferred total ordering before one can define robustness. This is clearly advantageous in the settings where a preference exists on the outset. But in general, as with the RTS approach, one typically has to keep track of multiple preferences at the beginning before refining them later on. Alternatively, as described earlier in Section 3.1, one can also view the STR approach as a computational strategy that is posed in order to solve some set-valued optimisation problem (8).
@@ 3.4.2 Commutative operations
3.4.3 Relation with reinforcement learning
@@ 4 Robust Pareto fronts
The Pareto front is an important quantity in standard multi-objective optimisation (2). It gives us valuable information about the different trade-offs that are happening in the objective space. In practice, this knowledge is typically used by a decision maker in order to help them make an informed and effective decision. So far, much of the work on robust multi-objective optimisation has focussed on classifying and identifying inputs that are deemed as robust in some sense. Very little work has focussed on the equally important problem of interpreting the robust trade-off that is happening in the objective space. This section addresses this gap in the literature by proposing a principled way to define a robust Pareto front under the RTS and STR perspectives for the length scalarised problems. More precisely, our work leverages ideas from a recent paper by Tu et al. (2024b), who showed how it is possible to reconstruct an interpolated Pareto front surface from a collection of length scalarised values. This key result was referred to as the polar parametrisation of a Pareto front surface, which we first recall in Section 4.1. Then we move on to Section 4.2 where we exploit this parametrisation result in order to define the general concept of a risk statistic associated with a random Pareto front surface (Proposition 4.1). Afterwards, in Section 4.3, we present the novel concept of an RTS front (Definition 4.3) and a STR front (Definition 4.4). Formally, these are the surfaces described by the corresponding RTS and STR length scalarised values, respectively. As we prove in the main results in this section, Proposition 4.2 and Proposition 4.3, both of these surfaces are Pareto front surfaces, under mild assumptions, and therefore are valid robust generalisations to the standard Pareto front trade-off surface. Finally, as an aside, in we present a theoretical result in Section 4.4 regarding the ordering of the extreme case Pareto front surfaces.
4.1 Polar parametrisation
To begin with, we recall the definition of a Pareto front surface in Definition 4.1. Intuitively, this surface is the one obtained by interpolating the strict Pareto front of a set using the weak Pareto partial ordering. As this interpolation could extend indefinitely, we consider truncating this surface at some lower bound defined by a reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. Afterwards, we recall the definition of a polar surface in Definition 4.2. Geometrically, a polar surface is any set that can be obtained by stretching the positive space of unit vectors $\mathcal{S}^{M-1}_{+}$ along the radial directions, before translating the origin to the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. Consequently, this implies that a polar surface can be described entirely by its projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}$, which computes the amount of stretching that occurs along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. As shown by Tu et al. (2024b), and recalled in Theorem 4.1, every Pareto front surface is a polar surface. In particular, it is a polar surface whose projected length function can be written in terms of the length scalarisation function (5). On the left of Figure 4, we present a pictorial illustration of this parametrisation result for the Pareto front surface described in Figure 1.
@@ 4.1 Polar parametrisation
Theorem 4.1 (Polar parametrisation)
(Tu et al., 2024b, Theorem 3.1) For any bounded set of vectors $A\subset\mathbb{R}^{M}$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if the corresponding Pareto front surface is non-empty $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\neq\emptyset$, then it admits the following polar parametrisation:
@@ 4.2 Risk statistics
To apply the polar parametrisation result on the objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, with a reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, we need to ensure that it fulfils the necessary assumptions outlined in Theorem 4.1. That is, we need to ensure that the objective function is bounded and that the corresponding Pareto front surface is non-empty. In this work, we only assume that the objective function is bounded. As a result, the polar parametrisation of any bounded set of points could potentially degenerate. Note that this event can however be avoided if the reference vector is set more judiciously (Remark 4.1). Formally, in this work, we consider the family of polar surfaces
@@
Figure 4: We illustrate the polar parametrisation (Theorem 4.1) and the Pareto front surface statistics (Proposition 4.1) for a simple two-dimensional example. On the left, we plot the polar parametrisation of the Pareto front surface from Figure 1. On the remaining plots, we draw the corresponding risk statistics associated with a finite collection of Pareto front surfaces and the partially coherent univariate risk functionals listed in Section 2.3.1.
@@ 4.3 Robust Pareto front surfaces
The polar parametrisation result in Theorem 4.1 tells us that we can reconstruct a Pareto front surface from its corresponding collection of maximum length scalarised values. We will now take advantage of this result in order to define the corresponding polar surfaces for the RTS and STR length scalarised problems. More specifically, we define the RTS front (Definition 4.3) and STR front (Definition 4.4) as the polar surface obtained using the RTS length values (14) and the STR length values (16), respectively.
@@
In Proposition 4.2 and Proposition 4.3, we present the main results in this section which states that these polar surfaces (20) and (21) are indeed valid Pareto front surfaces. Not only that, we give an explicit representation on what these Pareto front surfaces actually are. In words, the RTS front is the Pareto front surface associated with the union robust problem (13). Whilst the STR front is the Pareto front surface associated with the risk statistic of each input’s Pareto front surface distribution. The proof of these result are presented in Section A.7 and Section A.8, respectively.
@@ 4.4 Bounding a nominal surface
For any bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$ and strongly dominated reference vector $\boldsymbol{\eta}\in\cap_{\mathbf{x}\in\mathbb{X}}\cap_{\mathbf{y}\in f(\mathbf{x},\Xi)}\mathbb{D}_{\prec\mathrel{\mkern-5.0mu}\prec}[\{\mathbf{y}\}]$, we can write the target Pareto front surface in its polar parametrised form (18). By definition, this implies that the projected lengths of this Pareto front surface $Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\xi}^{*})$, along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$, can be written as
@@
where the term $z^{(m)}_{\boldsymbol{\eta},\boldsymbol{\lambda}}(y)=(y-\eta^{(m)})/\lambda^{(m)}$ is used to denote the $m$-th argument in the length scalarisation function (5) for any number $y\in\mathbb{R}$. It can be easily shown that this Pareto front surface can be bounded from above and below by the extreme cases of the Pareto front surface statistics, RTS fronts and STR fronts. Mathematically, this statement can be formalised by the following three inequalities:
@@
which all hold simultaneously for every uncertain parameter $\boldsymbol{\xi}^{*}\in\Xi$. We see clearly that the only difference between these bounds are the ordering of the three main operations: maximisation over the inputs, robustification over the uncertain parameters and scalarisation of the objective. In words, Proposition 4.4 states that the ordering of these three operations can be exploited in order to get a smaller bounding interval. The proof of this result is presented in Section A.9.
@@
Note that in the special case when the set of uncertain parameters is a singleton set $\Xi=\{\boldsymbol{\xi}^{*}\}$, all of these extreme case Pareto front surfaces are equal. In other words, the difference between these three approaches only begins to transpire when the uncertainty set grows. Proposition 4.4 gives us an ordering on the range of Pareto front surfaces one can obtain. It does not however give us any information regarding which method is the most effective to identify or estimate this uncertain parameter $\boldsymbol{\xi}^{*}\in\Xi$. This latter problem turns out to be case specific as it relies heavily on the assumptions that are placed on the objective function and the form of the uncertainty.

### v2

@@ 2.3 Risk functionals
Most existing work in robust multi-objective optimisation has focussed on the simple scenario where each uncertain parameter $\boldsymbol{\xi}\in\Xi$ is given equal weight. In this work, we consider a more general treatment where the uncertain parameter is treated as a random variable that is defined sensibly over a standard probability space. Precisely, we suppose that the uncertain parameter $\boldsymbol{\xi}\in\Xi$ is distributed according to some probability density $p\in\mathcal{P}[\Xi]$ that could potentially be known but might in general be unknown. Equipped with this set-up, the value of an input $\mathbf{x}\in\mathbb{X}$ will no longer be determined by its equally weighted output set (7), but instead will be defined using a sensible weighted alternative. More specifically, we will consider the use of risk functionals (Rüschendorf, 2013, Part 2) in order to obtain these proxy set of values. For this reason, we will now give a brief overview on the concepts of a univariate risk functional (Section 2.3.1) and a multivariate risk functional (Section 2.3.2). In Figure 2, we give an illustration of some univariate and multivariate risk functionals for a simple one-dimensional and two-dimensional example, respectively.
@@ 2.3.1 Univariate risk functionals
In the following sections, we will consider the problem of maximising the output of a risk functional. We will often refer to the resulting output of a univariate risk functional as the risk-adjusted or risk-penalised objective value. Naturally, we want these risk functionals to satisfy some sensible properties in order for the corresponding risk-penalised optimisation problem to make sense from a modelling perspective. For this reason, we focus our attention on partially coherent univariate risk functional (Definition 2.4). By design, these functionals satisfy four additional properties which are sufficient enough to prove many of the results which we state and describe later in Section 4. Note that all of the univariate risk functionals listed above (Examples 2.2, 2.3, 2.4 and 2.5) are partially coherent.
@@ 2.3.2 Multivariate risk functionals
$\boldsymbol{\rho}^{\textnormal{WorstCase}}[h]$ :=$:=\{(\inf_{\boldsymbol{\xi}\in\mathcal{E}}h^{(1)}(\boldsymbol{\xi}),\dots,\inf_{\boldsymbol{\xi}\in\mathcal{E}}h^{(M)}(\boldsymbol{\xi}))\},$
$\boldsymbol{\rho}^{\textnormal{BestCase}}[h]$ :=$:=\{(\sup_{\boldsymbol{\xi}\in\mathcal{E}}h^{(1)}(\boldsymbol{\xi}),\dots,\sup_{\boldsymbol{\xi}\in\mathcal{E}}h^{(M)}(\boldsymbol{\xi}))\}.$
@@
Given a probability density $p\in\mathcal{P}[\Xi]$, the multivariate expectation functional maps any vector-valued function $h:\Xi\rightarrow\mathbb{R}^{M}$ to the singleton set comprised of its corresponding expectation $\boldsymbol{\rho}^{\textnormal{Exp}}[h]:=\{\mathbb{E}_{p(\boldsymbol{\xi})}[h(\boldsymbol{\xi})]\}$. For example, Deb and Gupta (2005) considered a special case where the uncertain parameter models any additive input noise arising in (2), that is $\boldsymbol{\rho}^{\textnormal{Exp}}[f(\mathbf{x},\cdot)]=\{\mathbb{E}_{p(\boldsymbol{\xi})}[g(\mathbf{x}+\boldsymbol{\xi})]\}$ for any input $\mathbf{x}\in\mathbb{X}$, where $p\in\mathcal{P}[\Xi]$ is the input noise distribution and $g:\mathbb{X}\rightarrow\mathbb{R}^{M}$ is the vector-valued objective function.
@@
$\boldsymbol{\rho}^{\textnormal{CW}}[h]:=\{(\rho^{(1)}[h^{(1)}],\dots,\rho^{(M)}[h^{(M)}])\},$
@@ 3 Robust multi-objective optimisation
Several notions of robustness have already been proposed for the robust multi-objective optimisation problem (1). Regardless of these different definitions, the most prominent solution strategies that have been proposed to solve these robust problems are all largely based on the scalarisation methodology. We now proceed to introduce the two main solution strategies that will be the primary focus of this work: the robustify then scalarise (8) approach and the scalarise then robustify (11) approach.
3.1 Robustify then scalarise
Union robust optimisation problem
Recall from Section 2.2 that it is possible to recover any weakly Pareto optimal solution by solving a collection of length scalarised problems—see Theorem 2.1. We will now present the corresponding analogue of this result for the RTS setting.
Proposition 3.1 (Union robust weak solutions)
To elaborate, eq. 9 shows that the corresponding RTS length scalarised problem (9) can be used in order to target all of the weak solutions to the bounded vector-valued optimisation problem
3.2 Scalarise then robustify
STR length scalarised problem
3.3 Discussion
The RTS and STR approach are two general computationally-driven strategies for robust multi-objective optimisation. These two approaches are very flexible and lightweight because they only require the decision maker to specify a sensible family of scalarisation functions and a risk functional in order to define robustness. The former set of functions is determined by the decision maker’s internal preferences and specific goals. Whilst the latter functional is determined by the uncertainties in the problem and the decision maker’s adversity to risk. In this section, we will first discuss the relationships between our methodology and earlier work (Section 3.3.1); we will then look at some natural questions that one might have regarding these two approaches, such as: which approach should one use in practice (Section 3.3.2) and when are these two approaches equivalent (Section 3.3.3)? Moreover, as an aside, we also relate some of these ideas with existing concepts from multi-objective reinforcement learning (Section 3.3.4).
3.3.1 Related work
@@
In general, we cannot solve the set-valued optimisation problem (13) in its entirety. Instead, most existing optimisation strategies work by identifying some finite approximation to this problem based on some sensible adaptation to the scalarisation approach (Section 2.2). To elaborate, one usually aims to define and solve a finite collection of scalar-valued optimisation problems whose solutions are robust with respect to the set partial ordering of interest. Below, we will recall some of these examples for the upper and lower set ordered problem. Moreover, we highlight how these examples are just special cases of our RTS and STR approach.
@@ 3.3.2 When to robustify and scalarise?
In practice, decision makers are interested in making principled decisions that can be justified and explained to other stakeholders. From an interpretation perspective, the RTS approach is appealing because it can easily be interpreted as a standard multi-objective optimisation problem (2) wherein one replaces the objective function with a risk-penalised version (10). The corresponding scalarisation methodology can then be viewed as an operational technique in order to acquire the robust points which are the most desirable. In contrast, the STR approach takes a much more preference-driven strategy that is based largely on the scalarisation methodology. Specifically, one has to first identify a scalarisation function that describes the most preferred total ordering before one can define robustness. This is clearly advantageous in the settings where a preference exists on the outset. But in general, as with the RTS approach, one typically has to keep track of multiple preferences at the beginning before refining them later on. Alternatively, as described earlier in Section 3.3.1, one can also view the STR approach as a computational strategy that is posed in order to solve some set-valued optimisation problem (13).
@@ 3.3.3 Commutative operations
3.3.4 Connections with reinforcement learning
@@ 4 Robust Pareto fronts
The Pareto front is an important quantity in standard multi-objective optimisation (2). It gives us valuable information about the different trade-offs that are happening in the objective space. In practice, this knowledge is typically used by a decision maker in order to help them make an informed and effective decision. So far, much of the work on robust multi-objective optimisation has focussed on classifying and identifying inputs that are deemed as robust in some sense. To the best of our knowledge no work has focussed on identifying and interpreting the robust trade-off that is happening in the objective space. This section addresses this gap in the literature by proposing a principled way to define a robust Pareto front under the RTS and STR perspectives for the length scalarised problems. More precisely, our work leverages ideas from a recent paper by Tu et al. (2024b), who showed how it is possible to reconstruct an interpolated Pareto front surface from a collection of length scalarised values. This key result was referred to as the polar parameterisation of a Pareto front surface, which we first recall in Section 4.1. Then we move on to Section 4.2 where we exploit this parameterisation result in order to define the general concept of a risk statistic associated with a random Pareto front surface (Proposition 4.1). Afterwards, in Section 4.3, we present the novel concept of an RTS front (Definition 4.3) and a STR front (Definition 4.4). Formally, these are the surfaces described by the corresponding RTS and STR length scalarised values, respectively. As we prove in the main results in this section, Proposition 4.2 and Proposition 4.3, both of these surfaces are Pareto front surfaces, under mild assumptions, and therefore are valid robust generalisations to the standard Pareto front trade-off surface. Finally, as an aside, in we present a theoretical result in Section 4.4 regarding the ordering of the extreme case Pareto front surfaces.
4.1 Polar parameterisation
To begin with, we recall the definition of a Pareto front surface. Intuitively, this surface is the one obtained by interpolating the strict Pareto front of a set using the weak Pareto partial ordering. As this interpolation could extend indefinitely, we consider truncating this surface at some lower bound defined by a reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. Afterwards, we recall the definition of a polar surface.
@@ 4.1 Polar parameterisation
Geometrically, a polar surface is any set that can be obtained by stretching the positive space of unit vectors $\mathcal{S}^{M-1}_{+}$ along the radial directions, before translating the origin to the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. Consequently, this implies that a polar surface can be described entirely by its projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}$, which computes the amount of stretching that occurs along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. As shown by Tu et al. (2024b), and recalled in Theorem 4.1 below, every Pareto front surface is a polar surface. In particular, it is a polar surface whose projected length function can be written in terms of the length scalarisation function (5). On the left of Figure 4, we present a pictorial illustration of this parameterisation result for the Pareto front surface described in Figure 1.
Theorem 4.1 (Polar parameterisation)
(Tu et al., 2024b, Theorem 3.1) For any bounded set of vectors $A\subset\mathbb{R}^{M}$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if the corresponding Pareto front surface is non-empty $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\neq\emptyset$, then it admits the following polar parameterisation:
@@ 4.2 Risk statistics
To apply the polar parameterisation result on the objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, with a reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, we need to ensure that it fulfils the necessary assumptions outlined in Theorem 4.1. That is, we need to ensure that the objective function is bounded and that the corresponding Pareto front surface is non-empty. In this work, we only assume that the objective function is bounded. As a result, the polar parameterisation of any bounded set of points could potentially degenerate. Note that this event can however be avoided if the reference vector is set more judiciously (Remark 4.1). Formally, in this work, we consider the family of polar surfaces
@@
Figure 4: We illustrate the polar parameterisation (Theorem 4.1) and the Pareto front surface statistics (Proposition 4.1) for a simple two-dimensional example. On the left, we plot the polar parameterisation of the Pareto front surface from Figure 1. On the remaining plots, we draw the corresponding risk statistics associated with a finite collection of Pareto front surfaces and the partially coherent univariate risk functionals listed in Section 2.3.1.
@@ 4.3 Robust Pareto front surfaces
The polar parameterisation result in Theorem 4.1 tells us that we can reconstruct a Pareto front surface from its corresponding collection of maximum length scalarised values. We will now take advantage of this result in order to define the corresponding polar surfaces for the RTS and STR length scalarised problems. More specifically, we define the RTS front and STR front as the polar surface obtained using the RTS length values (9) and the STR length values (12), respectively.
@@
In Proposition 4.2 and Proposition 4.3, we present the main results in this section which states that these polar surfaces (20) and (21) are indeed valid Pareto front surfaces. Also, we give an explicit representation on what these Pareto front surfaces actually are. In words, the RTS front is the Pareto front surface associated with the union robust problem (10). Whilst the STR front is the Pareto front surface associated with the risk statistic of each input’s Pareto front surface distribution. The proof of these result are presented in Section A.7 and Section A.8, respectively.
@@ 4.4 Bounding a nominal surface
For any bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$ and strongly dominated reference vector $\boldsymbol{\eta}\in\cap_{\mathbf{x}\in\mathbb{X}}\cap_{\mathbf{y}\in f(\mathbf{x},\Xi)}\mathbb{D}_{\prec\mathrel{\mkern-5.0mu}\prec}[\{\mathbf{y}\}]$, we can write the target Pareto front surface in its polar parameterised form (18). By definition, this implies that the projected lengths of this Pareto front surface $Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\xi}^{*})$, along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$, can be written as
@@
where the term $z^{(m)}_{\boldsymbol{\eta},\boldsymbol{\lambda}}(y)=(y-\eta^{(m)})/\lambda^{(m)}$ is used to denote the $m$-th argument in the length scalarisation function (5) for any number $y\in\mathbb{R}$. It can be easily shown that this Pareto front surface can be bounded from above and below by the extreme cases of the Pareto front surface statistics (denoted by PS), RTS fronts and STR fronts. Mathematically, this statement can be formalised by the following three inequalities:
@@
which all hold simultaneously for every uncertain parameter $\boldsymbol{\xi}^{*}\in\Xi$. We see clearly that the only difference between these bounds are the ordering of the three main operations: maximisation over the inputs, robustification over the uncertain parameters and scalarisation of the objective.
@@
Proposition 4.4 states that the ordering of these three operations can be exploited in order to get a smaller bounding interval. The proof of this result is presented in Section A.9. Note that in the special case when the set of uncertain parameters is a singleton set $\Xi=\{\boldsymbol{\xi}^{*}\}$, all of these extreme case Pareto front surfaces are equal. In other words, the difference between these three approaches only begins to transpire when the uncertainty set grows. Proposition 4.4 gives us an ordering on the range of Pareto front surfaces one can obtain. It does not however give us any information regarding which method is the most effective to identify or estimate this uncertain parameter $\boldsymbol{\xi}^{*}\in\Xi$. This latter problem turns out to be case specific as it relies heavily on the assumptions that are placed on the objective function and the form of the uncertainty.